# A Comprehensive Survey on Automatic Text Summarization with Exploration of LLM-Based Methods

### v2

@@ 8 Applications based on ATS
The ultimate objective of Automated Text Summarization (ATS) is to enhance the efficiency of information retrieval and analysis in real-world scenarios. ATS has applications in various domains, including the summarization of news articles, scientific papers and other areas that involve substantial reading efforts.
News Summarization: represents the most extensively investigated domain within ATS, largely owing to the richness and maturity of available datasets, coupled with the significant demand for its application across various industries. [300] introduced sentence fusion, a text-to-text generation technique that merges phrases containing similar information into a single sentence for summary generation. Due to the real-time nature of news, [301] presented Newsblaster, a system that is based a clutering meethods for crawled news articles by topic, and thus produce summaries based on topic clusters. In addition, based on the chronological nature of the forward and backward relations of news articles, [302] proposes a time-line based summarization method,
@@ 9 Future Directions in Large Language Model-Based Summarization
The integration of large language models (LLMs) into automatic text summarization (ATS) represents a significant advancement in the field, enabling models to generate more coherent, contextually aware, and semantically accurate summaries. As LLMs continue to evolve, several promising future directions emerge:
Scaling and Efficiency: While LLMs, such as GPT and LLaMa-based models, have shown superior performance in text summarization, their computational cost remains a challenge. Future research could focus on optimizing the efficiency of these models, exploring techniques such as model distillation, pruning, and quantization to reduce computational overhead without sacrificing performance.
Domain Adaptation and Generalization: LLMs are typically trained on large, general-purpose corpora, which might not always translate well to domain-specific tasks. Future work can explore more robust domain adaptation methods, such as fine-tuning LLMs on specialized corpora or developing multi-task learning frameworks that enable LLMs to generalize across diverse domains, from medical texts to legal documents.
Explainability and Interpretability: The black-box nature of LLMs poses challenges for interpretability. Future research should aim at making LLM-based summarization methods more transparent, by developing techniques that explain how certain sentences or sections were prioritized in generating a summary. This is especially important in high-stakes domains, such as legal or financial summarization, where interpretability is critical.
Handling Longer Documents: Despite their strengths, LLMs still face difficulties when handling extremely long input texts, often truncating or missing important information. Research should focus on improving memory mechanisms or developing hierarchical approaches that allow LLMs to better capture and summarize longer documents, possibly by segmenting input and summarizing it in parts.
Multimodal Summarization: Given the success of LLMs in text processing, an exciting avenue for future work is expanding these models to handle multimodal summarization, where the input includes not only text but also images, audio, or video. LLMs, coupled with multimodal learning techniques, could provide richer, more comprehensive summaries from diverse input sources.
Ethical Considerations and Bias: As LLM-based methods are increasingly adopted, it becomes crucial to address issues of bias and fairness. Future research should focus on developing techniques to mitigate the propagation of biases in summaries, ensuring that outputs are fair and representative of diverse perspectives, especially when summarizing content from sensitive or polarizing domains.
Interactive and Personalized Summarization: The future may also see a shift towards more interactive or personalized summarization, where users can guide or customize summaries based on specific needs or preferences. Leveraging LLMs to incorporate user feedback in real-time or adapting summaries to individual styles and requirements could significantly enhance the utility of summarization systems.
By addressing these challenges and exploring these directions, LLM-based methods will continue to push the boundaries of automatic text summarization, offering more flexible, accurate, and scalable solutions.

### v3

@@ 9 Applications based on ATS
The ultimate objective of Automated Text Summarization (ATS) is to enhance the efficiency of information retrieval and analysis in real-world scenarios. ATS has applications in various domains, including the summarization of news articles, scientific papers, and other areas that involve substantial reading efforts.
News Summarization: represents the most extensively investigated domain within ATS, largely owing to the richness and maturity of available datasets, coupled with the significant demand for its application across various industries. [333] introduced sentence fusion, a text-to-text generation technique that merges phrases containing similar information into a single sentence for summary generation. Due to the real-time nature of news, [334] presented Newsblaster, a system that is based a clustering methods for crawled news articles by topic, and thus produces summaries based on topic clusters. In addition, based on the chronological nature of the forward and backward relations of news articles, [335] proposes a timeline-based summarization method,
@@ 10 Future Directions in Large Language Model-Based Summarization
Despite the remarkable progress of Large Language Models (LLMs) in Automatic Text Summarization (ATS), several limitations remain in current research. These include challenges in adapting general-purpose LLMs to domain-specific summarization tasks, handling long documents under the token length constraints of LLMs, and improving the interpretability of internal LLM mechanisms. In this section, we outline key future research directions in LLM-based ATS, grounded in the limitations identified throughout this survey.
Mitigating Hallucination and Enhancing Trust in LLMs for ATS: One of the key challenges in applying Large Language Models (LLMs), not only in Automatic Text Summarization (ATS) but across a wide range of NLP tasks, is the issue of hallucination, where the generated content is factually incorrect or not grounded in the source text. In the context of ATS, this results in summaries that may include fabricated details, misrepresentations, or unsupported claims, thereby compromising the reliability and trustworthiness of the output[349, 350, 351]. This problem is particularly pronounced in abstractive summarization, where LLMs generate novel sentences without strict reliance on the input content.
To address hallucination in ATS, future research could potentially explore three key directions. First, hallucination detection and evaluation require significant advancement. Conventional metrics such as ROUGE and BLEU rely heavily on lexical overlap and are inadequate for identifying factual inconsistencies. While newer metrics like BERTScore provide semantic similarity measures, they still fall short in assessing factual alignment. There is a pressing need for factuality-aware evaluation metrics that can reliably quantify the trustworthiness and source alignment of generated summaries. Second, grounded and retrieval-augmented summarization offers a potential direction to mitigate hallucination by conditioning generation on sentence-level evidence or extracted facts from the input. Techniques such as Retrieval-Augmented Generation (RAG) or hybrid extractive-abstractive frameworks can serve to anchor LLM outputs in verifiable content. Third, emerging work on reinforcement learning for factual alignment presents a compelling direction, where trust-oriented reward signals, based on external fact-checkers or entailment models, can guide LLMs toward producing summaries that are not only fluent but also faithful to the source.
Scalable Long-Document Summarization with LLMs: A critical limitation of current Large Language Model (LLM)-based Automatic Text Summarization (ATS) systems is their restricted ability to handle long-form documents due to input token constraints. Popular models such as GPT-3 and PaLM typically support context windows ranging from 2K to 32K tokens, which significantly limits their applicability to documents like legal rulings, scientific publications, or technical manuals that require reasoning across extended discourse structures [352, 353]. When input texts exceed these boundaries, truncation can lead to the loss of essential information, while naive chunking strategies often disrupt semantic coherence and degrade summary quality.
To address these challenges, one promising research direction is the development of long-context modeling strategies. These include memory-augmented transformers, recurrence-based architectures, and retrieval-augmented generation frameworks, which aim to handle longer sequences efficiently without linear increases in computational complexity [354, 355]. In parallel, chunk-wise summarization frameworks, which segment documents into coherent units and incorporate cross-segment reasoning or hierarchical attention mechanisms, can help maintain contextual continuity and improve overall summary cohesion. Exploring these directions will be essential for advancing the scalability and robustness of LLM-based summarization systems in real-world, document-intensive scenarios.
Improving Interpretability in LLM-Based ATS: Despite the fluency and human-likeness of summaries generated by LLM-based ATS systems, the underlying generation process often remains a black box, offering limited transparency into how specific content is derived from the source text. This lack of interpretability poses a critical limitation, particularly in applications where accuracy and accountability are essential. Users may hesitate to trust or adopt ATS systems unless they can clearly verify the provenance, contextual relevance, and justification of each piece of generated information.
To improve interpretability in LLM-based ATS, future research should explore two potential directions grounded in recent advancements in large language model techniques. First, prompt-based interpretability, including Chain-of-Thought (CoT) prompting and structured prompt engineering, offers a lightweight yet effective way to make the generation process more transparent. By encouraging models to generate intermediate reasoning steps or explicitly cite relevant portions of the input, these approaches help users understand the logical flow behind the summary content without requiring model retraining. Second, retrieval-augmented summarization (RAG) can improve interpretability by linking generated summaries to specific, verifiable source passages retrieved at inference time. These directions collectively move toward ATS systems that not only produce high-quality summaries but also provide clear, verifiable reasoning aligned with user expectations and real-world accountability standards.
Domain-Specific Summarization with LLMs: While Large Language Models (LLMs) have demonstrated good capabilities in generic summarization tasks, their performance often degrades in domain-specific scenarios such as biomedical literature, legal documents, financial reports, or technical manuals. These domains present unique challenges, including specialized terminology, complex discourse structures, and high factual precision requirements, which general-purpose models trained on open-domain corpora are inappropriate to handle. As a result, LLM-generated summaries in these settings may lack accuracy, omit critical content, or fail to meet the precision and clarity expected by domain experts.
To address these challenges, future research could potentially explore three directions. First, domain-adaptive fine-tuning remains a widely used and effective approach, where LLMs are further trained on in-domain corpora to internalize specialized knowledge. Notable examples based on the pre-trained models include BioBERT for biomedical texts [356], Legal-BERT for legal documents [357], and SciBERT for scientific articles [358]. To reduce the computational burden of full model retraining, parameter-efficient tuning methods such as LoRA [257] offer a promising alternative by enabling lightweight adaptation with minimal additional parameters. Second, prompt-based adaptation strategies, such as in-context learning or few-shot prompting using domain-specific exemplars, can guide LLMs toward stylistically appropriate and contextually relevant summaries without modifying model weights [359]. Third, integrating domain knowledge from structured resources (e.g., ontologies, taxonomies, or controlled vocabularies) through retrieval-augmented or knowledge-grounded generation frameworks [360] can further enhance the factuality and informativeness of the output. These directions collectively aim to build robust summarization systems that are better aligned with the demands of specialized domains and capable of producing trustworthy, high-utility summaries for professional use.
@@ 11 Conclusion
Appendix A Annotation Questionnaire Design
A.1 A.1 is_LLM Classification Task
Question: Determine whether the paper is related to Large Language Models (LLMs) based on its title, abstract, and main content. If “Yes”, provide a brief rationale; if “No”, explain the unrelated domain.
Judgment Criteria:
  1. 0.
    Yes: Core content involves LLM architecture design, training methods, performance optimization, or application scenarios (e.g., GPT, BERT, Transformer).
  2. 1.
    No: Focuses on traditional machine learning, computer vision, small-scale NLP models (e.g., RNN, CNN), or non-model-related fields (e.g., theoretical analysis, hardware acceleration).
Examples:
  • 1.
    Yes: The paper proposes an improved LLaMA-3 architecture.
  • 2.
    No: The paper studies traditional SVM applications in ATS.
A.2 A.2 is_dataset Classification Task
Question: Judge whether the paper is related to dataset construction. If “Yes”, describe the dataset’s purpose and characteristics; if “No”, explain the reason.
Judgment Criteria:
  1. 0.
    Yes: The paper explicitly constructs and releases a new dataset or details the extension/improvement of an existing dataset (e.g., annotation methods, data sources).
  2. 1.
    No: The paper only uses publicly available datasets without involving dataset design or improvement.
Examples: - Yes: The paper introduces the VSoLSCSum dataset with annotation tools and benchmarks. - No: The paper employs the Multi-News dataset for training but does not discuss dataset design.
A.3 A.3 Type Classification Task
Question: Select the paper’s task type from the following categories and provide justification:
  1. 0.
    Evaluation: Systematic comparison of model/method performance.
  2. 1.
    Survey: Survey of summarization research progress, challenges, or applications.
  3. 2.
    Model: Proposes a novel architecture or improves an existing model.
  4. 3.
    Others: Other types, please indicate the specific category.
Examples:
  • 1.
    Evaluation: The paper evaluates GPT-3’s summarization performance via human preference judgments.
  • 2.
    Model: The paper proposes a sparse attention mechanism for LLMs.
A.4 A.4 Methodology Classification Task
Question: Select the technique(s) involved in the paper (multiple choices allowed) and briefly describe their application:
  1. 0.
    Prompt Engineering: Use prompts to guide model outputs.
  2. 1.
    Fine-tuning: Train pre-trained models on specific tasks.
  3. 2.
    Distillation: Compress large models into smaller ones via knowledge distillation.
  4. 3.
    Others: Specify the method.
Examples:
  • 1.
    Prompt Engineering: The paper designs a prompt template to enhance model generalization.
  • 2.
    Distillation: The paper distills GPT-3’s knowledge into a smaller T5 model.
A.5 A.5 Domain Classification Task
Question: Classify the paper’s domain, selecting from the following categories: News, Novel, Scientific Paper, Blog, Dialogue, Medical, or Other. Provide a rationale based on the title, abstract, and main content.
Classification Criteria:
  1. 0.
    News: Summarization of journalistic articles or current events.
  2. 1.
    Novel: Summarization of fictional literary works.
  3. 2.
    Scientific Paper: Summarization of academic research articles.
  4. 3.
    Blog: Summarization of personal or professional blog posts.
  5. 4.
    Dialogue: Summarization of conversation-based texts (e.g., interviews, meetings).
  6. 5.
    Medical: Summarization of clinical notes, medical reports, or health-related content.
  7. 6.
    Other: Specify the domain if not listed above.
Examples:
  • 1.
    Scientific Paper: The paper presents a model for arxiv papers summarization.
  • 2.
    Medical: The study proposes a model for condensing discharge summaries in electronic health records.
A.6 Annotation Guidelines
  • 1.
    Base judgments on the paper’s title and abstract.
  • 2.
    Document controversies in the “Remarks” section if ambiguities arise.
  • 3.
    Preserve original rationales (e.g., cite key sentences from the abstract).
Appendix B Author Biographies
Yang Zhang
Yang Zhang is an assistant professor with the Southwestern University of Finance and Economics, China. He received his Ph.D. from the Graduate School of Informatics, Kyoto University in 2022. Before that, he received his Bachelor of Engineering from the University of New South Wales and his Master of Economics and Finance from Sydney University. He joined Southwestern University of Finance and Economics, China in 2023 as a lecturer. His research interests include text mining, text recommendation, and NLP for financial technology.
Hanlei Jin
Hanlei Jin is currently working toward the Ph.D. in Management Science and Engineering at Southwestern University of Finance and Economics, Chengdu, China. He holds a B.S. degree in Management from the same institution. His research interests include text mining, text generation, and financial intelligence.
Dan Meng
Dan Meng is a professor with the Southwestern University of Finance and Economics, Chengdu, China. Previously, she received her Ph.D. from Southwest Jiaotong University, Chengdu, China. Her research interests include intelligent finance, intelligent decision making, and uncertainty information processing.
Jun Wang
Jun Wang is a professor with the Southwestern University of Finance and Economics, China. Prior to his current appointment, he was a researcher at the Memorial University of Newfoundland in St. John’s, Canada, and he was awarded the National Scholarship in 2017. His research interests include NLP, social media, social network analysis, financial analysis, and business intelligence.
Jinghua Tan
(Member, IEEE) Jinghua Tan received the B.S., M.S., and Ph.D. degrees from the Southwestern University of Finance and Economics, Chengdu, China. She is an associate professor with Sichuan Agricultural University and was a visiting scholar at the Memorial University of Newfoundland in St. John’s, Canada in 2019. Her research interests include data mining and financial intelligence.